Circular No. 8557

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 N1 (JUELS-HOLVORCEM) C. W. Juels, Fountain Hills, AZ; and P. R. Holvorcem, Campinas,

Brazil, report their discovery of a diffuse comet on CCD images taken with a 0.07-m f/2.8 refractor at Fountain Hills (discovery observation given below). Following posting on the ‘NEO Confirmation Page’, other CCD observers have remarked on the object’s cometary nature. J. E. McGaha (Tucson, AZ, 0.36-m f/10 Schmidt-Cassegrain reflector) notes that ten stacked 60-s exposures taken on July 3.4 UT show a bright inner coma of diameter 30′′ and a faint outer coma extending out to a diameter of 70′′, with a 20′′-wide tail that is 120′′ long in p.a. 255^{o}. C. W. Hergenrother, Lunar and Planetary Laboratory, writes that a 300-s R-band exposure with the University of Arizona 1.54-m reflector shows a circular coma 27′′ in diameter, but with no hint of a tail. Images taken by L. Buzzi and F. Luppi (Varese, Italy, 0.60-m f/4.6 reflector) on July 3.1 show a diffuse, round coma of diameter ∼ 20′′. The available astrometry (with prediscovery observations by Juels and Holvorcem on June 30 and July 1), preliminary parabolic orbital elements [T = 2005 Aug. 21.24 TT, ω = 78^{o}.94, Ω = 3^{o}.04, i = 50^{o}.85 (equinox 2000.0), q = 1.1380 AU], and an ephemeris appear on MPEC 2005-N10.

2005 UT  α_{2000}  δ_{2000}  Mag.
July 2.44762  3^{h}16^{m}42.^{s}45  +34^{o}57′46′′.1  14.6

SUPERNOVAE 2005cp AND 2005ct Further to IAUC 8550, J. Graham and W. Li report the LOSS discovery

of an apparent supernova (mag 17.3) on unfiltered KAIT images taken on July 1.49 and 2.47 UT. SN 2005ct is located at α = 0^{h}39^{m}40^{s}.50, δ = −14^{o}14′16′′.6 (equinox 2000.0), which is 3′′.0 west and 2′′.3 south of the center of NGC 207. A KAIT image taken on Jan. 20.19 showed nothing at this position (limiting mag ∼ 19.5).

M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center for Astrophysics, report that a spectrogram (range 340–740 nm) of SN 2005cp (cf. IAUC 8550), obtained by R. Hutchins on June 30.43 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-IIn supernova. The spectrum consists of a blue continuum, indicating young age, and Balmer emission lines with a narrow component of ∼ 400 km/s (FWHM) superposed on a weaker and broader component of ∼ 4000 km/s (FWHM), after removing the NED recession velocity of 6616 km/s for the host galaxy (from Falco et al. 1999, PASP 111, 438).

2005 July 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green